Using the data below, create a time-phased budget for the project. Show how much the daily and cumulative costs for the project are, just as the monthly and cumulative costs are shown in Exhibit 10.9.

Immediate Duration Hourly Hours per Week WBS Activity Predecessor in Weeks Resource Cost Operational definition Research literature Becky 1.1 3 $35 30 1.2 Identify and define terms Ann $60 1.1 1 30 $50 1.3 Obtain approval of definition Clive 1.2 2 Target Selection 2 Solicit partners for pilot Ann 2.1 2 Hold brainstorming meeting Becky 2,2 2.1 2 Identify characteristics of targets 2.2, 3.1 Ann 2,3 1 Obtain approval of partners Clive 2,4 2.3, 1.2, 3.4 1 Question set Identify process group members Clive 3.1 2 Develop question set 4 3,2 2.3 Ann 3.3 Prototype and validate question set Becky 3.2 3 Add partners Becky 3.4 3.1, 2.1 3 20
Duration Hours per Immediate Hourly WBS Predecessor in Weeks Resource Cost Week Activity Pilot process 4 Schedule with target audience Becky 4.1 2,4 Conduct beta test Clive 4.2 3.4, 2.4 2 Process feedback from target 4.3 4.2 2 Ann audience Conduct pilot Clive 4.4 4.3 2 Analyze results Clive 4.5 4.4 2
EXHIBIT 10.9 AGGREGATION OF PROJECT BUDGET FEB MAR APR MAY JUN WP 1221 WP 1222 WP 1231 WP 1232 WP 1233 WP 1241 WP 1221 ($10,000)$10,000 WP 1222 ($10,000) WP 1231 ($45,000) WP 1232 ($20,000) WP 1233 ($15,000) WP 1241 ($20,000) $5,000 $5,000 $15,000 $30,000 $20,000 $15,000 $10,000 $10,000 Cost per month $15,000 $20,000 $50,000 $25,000 $10,000 Cumulative cost $15,000 $35,000 $85,000 $110,000$120,000 $120K $100K $80K $60K $40K $20K 0 Source: Kevin P. Grant, University of Texas, San Antonio. Adapted with permission.
Answer #1

Time phased budget and Cumulative cost chart are constructed as below:

Cost per day for Becky = 35*30/5 = 210

Cost per day Ann = 60*30/5 = 360

Cost per day for Clive = 50*20/5 = 200

Time Phased Budget for the project Day Day Day Day Day Day Day Day Day Day Day Day Day Day WBS Predecessor Duration Resource